Briefly answer

1). In an FE simulation with Abaqus, if the unit for the length is μm and the unit for the force is Netwon, what is the correct unit for the elastic modulus? Why?

11). What is a nonlinear elastic material?

12). The ratio of the initial bulk modulus to the shear modulus is 10,000 for a rubber. What is the rubber material’s Poisson’s ratio?

13). In a frictionless elastic contact between two cylinders with parallel axis, where does the maximum shear stress occur?

14). List two approaches used by Abaqus to simulate surface-based contact.

15). What is the option to use finite deformation in Abaqus/Standard?
Answer #1

Solution ie everything In Abaqus, there is no unit definition for Input and Output parameters is relative. The unit of output3. K G = 10,000 we know E = 2G (144) E = 3K (1-24). * (sk (1.8)) = f(26 (1441)) 3 (1-90) (1424) X G 10,000 x 9 (1-94) = (1404followed are s. General approaches that are (a) General contact algorithm which allows very simple definitions of contact witthank you.